Fermi National Laboratory
 

Accelerator Update for March 26, 2007 - March 28, 2007

 

Wednesday March 28

The day shift began with Operations monitoring Tevatron (TeV) store 5301 (48.54E30) the Recycler stash (224.7E10), with the Antiproton Source (Pbar) stacking (8.8E10), with the Booster, Linac, Main Injector (MI), and Recycler operating normally, with MiniBooNE off, with NuMI taking beam, and with D-Zero (D0) and the Collider Detector at Fermilab (CDF) using the colliding beams.

At 8:21 AM, MI aborted due to the MI-8 collimators. It reset.

At 8:26 AM, the NuMI beam permit tripped for the first of two times due to beam positions. It reset.

At 9:42 AM, MI tripped off again due to the MI-8 collimators.

A Pbar expert began a target sweeping study at 11:22 AM. The study lasted about two hours.

At 3:47 PM, Operations reported that they had lost communications with the Meson MS2 area. Operators investigated.

Operations terminated store 5301 at 4 PM.

Operations reported at 4:11 PM that the Duty Electrician had restored power to the Meson MS2 area.

Operations reported at 4:14 PM that the Recycler had stashed antiprotons twice during the day shift.

At 4:48 PM, LRF3 tripped off. It reset.

At 5:47 PM, the TeV aborted. Operations investigated and found that two CDF CAMAC crates were not responding.

At 6:17 PM, LRF3 tripped off. It reset.

At 6:28 PM, Operators reported that there was no power for the two CDF CAMAC crates. Operators continued to investigate.

At 7:03 PM, MI tripped off due to the MI-8 collimators for the first of three times in two minutes. It reset each time.

Operations reported 7:13 PM that the CDF CAMAC crates, the high voltage for beam monitoring, and other equipment in the CDF control room were getting their power from a panel in the TeV tunnel and were all off.

Operations had prepared the TeV for access earlier and at 8:55 PM reported that the TeV tunnel was dark and the A50 panel was dead. Operators traced the power to two panels next to each other by the A and B sector gates. These 120-volt panels were also down and one of them was wet. The operators continued to trace the power to a panel in the B0 compressor room: a 480V breaker. The Duty Electrician couldn’t reset the breaker. At 9:21 PM, he reported that water was running through the second panel by the A sector gate and that was the problem.

The Duty Electricians began the TeV panel repair at 10:48 PM.

Operations reported at 11:58 PM that the Recycler had stashed twice during the evening shift.

Thursday March 29

The midnight shift began with the TeV in access to repair a 120-volt AC panel, with Operations monitoring the stash (361.7E10), with Pbar stacking (39.6E10), and with NuMI taking beam.

At 1:12 AM, the H- Source tripped off due to a vacuum gauge problem. Operators accessed the Preaccelerator H- dome to power cycle the vacuum gauge and reset the VME. Linac resumed sending beam at 1:53 AM.

At 4:32 AM, the Duty Electricians reported that after extensive drying, the top breakers on the one panel were still bad and would have to be changed. They are looking for parts.

Ground water caused this problem. The picture on the left shows the location of the two panels. The picture on the right shows the two panels and a transformer. The bad breakers are circled.

Operations reported at 7:59 AM that the Recycler had stashed three times during the midnight shift.

CDF reported at 8:20 AM that they needed to access their collision hall to look for a possible helium leak. They had already looked everywhere else.

Operations reported that power returned to the TeV tunnel at B0 at 8:24 AM. Vacuum technicians accessed to check on vacuum pumps in the area.

A TeV expert reported at 8:52 AM that he had repaired the Tevatron Electron Lens (TEL1) that had tripped off.

CDF came out of access at 9:32 AM. The other TeV accesses had ended.

Operations began the TeV turn on at 10:09 AM.

Operations put the TeV through a dry squeeze at 12:13 PM.

At 12:25 PM, a sector D0 quadrupole (D0Q5) tripped off on a lead failure. It reset.

Operations prepared the Pbar PreVault area for an access at 12:59 PM to install some equipment.

Operations put the TeV through another dry squeeze at 1:19 PM.

Operations put the TeV through a 4x0 wet squeeze at 1:50 PM.

Operations put the TeV into shot setup at 2:24 PM.

Pbar came out of access. Operations established beam and an expert began tuning the beam at 3:33 PM.

Operations established store 5306 at 4:33 PM with an initial luminosity of 264.13E30.

The new Meson MTest T965 experiment reported at 5:09 PM that they were ready to take beam.

A Switchyard expert established beam to MTest at 5:20 PM.

MTest experimenters reported at 5:42 PM that they were through with beam for the night.

At 8:56 PM, the H- Source tripped off for the first of two times in about an hour. Operators accessed the dome after the second trip to power cycle the vacuum gauge.

Operations reported at 11:58 PM that the Recycler had stashed twice during the evening shift.

Friday March 30

The midnight shift began with Operations monitoring store 5306 (102.79E30) and the stash (172.2E10), with Pbar stacking (33.5E10), and with NuMI taking beam.

At 12:55 AM, LRF3 tripped off for the first of three times in an hour. It reset each time.

At 1:50 AM, the H- Source vacuum gauge tripped off. It reset.

At 4:03 AM, a vacuum problem tripped off the H- Source. Operators accessed the dome to power cycle the vacuum gauge.

At 7:07 AM, the TeV suffered a sector F4 quench due to a trim power supply (T:VA11) current sag. Operations called in a controls expert.

Operations made D0 ready for an access at 8:09 AM.

Operations reported at 8:16 AM that the Recycler had stashed three times during the midnight shift.

The Plans for Friday and the Weekend

The plans are to first complete the work on TeV power supply controls for VA11 and then wait for D0 to complete their access to repair a power supply for their forward muon detector. The plans for the reset of the day and the weekend are to stack, stash, and store. The Run Coordinator reports that next week he is planning a 12-hour shutdown so that CDF can drain and refill their cooling system. He will allow other machine work that doesn’t interfere with stacking and stashing.


Accelerator Update Archive

More Information
For Tevatron luminosity charts and the current status of Fermilab's accelerators and detectors (live!), please go to Fermilab Now.

Comments and Suggestions
What do you think about the Accelerator Updates? Please send comments and suggestions to: accelupdates@fnal.gov.

last modified 03/30/2007    email Fermilab